Use data from the table below to calculate the equilibrium constants at 25 ?C for each reaction. Standard Thermodynamic Quantities for Selected Substances at 25?C Substance ?H?f(kJ/mol) ?G?f(kJ/mol) S?(J/mol?K) ?G of PH3=13.5, ?G of H2=0, ?G of P4=24.4

3/2H2(g)+1/4P4(g)?PH3(g)
Answer #1

3/2H2(g) + 1/4P4(g) <----> PH3(g)

DG0rxn = (1*DG0f,PH3(g) ) - (3/2*DG0f,H2 + 1/4*DG0f,P4)

       = (1*13.5) - (3/2*0 + 1/4*24.4)

       = 7.4 Kj/mol

dG0 = -RTlnK

7.4*10^3 = -8.314*298lnk

k = equilibrium constant = 0.0504